How much energy is needed to melt a 100-g sample of aluminum whose initial temperature is 20 degrees celcius.




1
Expert's answer
2022-03-15T10:34:17-0400

Solution.

"m=0.1kg;"

"t_0=20^oC;"

"t=660^oC;"

"\lambda=390\sdot10^3J/kg;"

"c=920J\/{kg^oC};"

"Q=cm\Delta t+\lambda m;"

"Q=920\sdot0.1\sdot(660-20)+390\sdot10^3\sdot0.1=97880J;"

Answer: "Q=97880J."